Pramod goes to a shop to buy some clothes. Shopkeeper shows him 5 shirts, 6 pants and 13 t-shirts. If he selects three clothes at random, then find the probability that, randomly chosen clothes does not contain pant.

1
Expert's answer
2022-02-15T11:22:53-0500

There are "5+6+13=24" clothes. Pramod can select three clothes in "\dbinom{24}{3}=2024."

Then

"P(No\ pants)=\dfrac{\dbinom{5+13}{3}\dbinom{6}{0}}{\dbinom{24}{3}}=\dfrac{816(1)}{2024}"

"=\dfrac{102}{253}"

The probability that, randomly chosen clothes does not contain pant is "\dfrac{102}{253}\approx0.403."